Whirly corporation s contribution format
income statement for the most recent month is shown below:
                                             Total.            Per unit
Sales ( 7200 units)        223200. $         31.00 $
Variable expenses.        129600.           18.00
Contribution margin.      93600.             13.00
Fixed expenses.               55100
Net operating income.     38500
Required: ( consider each case independently)
1. What would be the revised net operating income per month if the sales volume increases by 70 units?
2. What would be the revised net operating income per month if the sales volume decreases by 70 units?
3. What would be the revised net operating income per month if the sales volume is 6200 units?
1 . Revised net operating income.
2 . Revised net operating income.
3 . Revised net operating income.
Answer #1
1 Sales volume=7200+70=7270 units
Revised net operating income:
Total Per unit
Sales ( 7270 units)         (7270*31) 225370 31
Variable expenses.     (7270*18) 130860 18
Contribution margin 94510 13
Fixed expenses 55100
Net operating income 39410
2 Sales volume=7200-70=7130 units
Revised net operating income:
Total Per unit
Sales ( 7130 units)         (7130*31) 221030 31
Variable expenses.     (7130*18) 128340 18
Contribution margin 92690 13
Fixed expenses 55100
Net operating income 37590
3 Sales volume=6200 units
Revised net operating income:
Total Per unit
Sales ( 6200 units)         (6200*31) 192200 31
Variable expenses.     (6200*18) 111600 18
Contribution margin 80600 13
Fixed expenses 55100
Net operating income 25500